OpenStudy (anonymous):

The volume of a cylinder is 81 cubic feet. What is the volume of a cone that fits exactly inside the cylinder?

OpenStudy (anonymous):

@Abmon98 Help with this one too? :P kinda the same

OpenStudy (abmon98):

sure :)

OpenStudy (anonymous):

Okay so what now :p

OpenStudy (hba):



OpenStudy (hba):

Now,Do you know the formula?

OpenStudy (anonymous):

The volume of a cone is 1/3 * pi * r^2 * H @hba

OpenStudy (anonymous):

@Abmon98

OpenStudy (abmon98):

is the radius or height given

OpenStudy (abmon98):

check your book

OpenStudy (anonymous):

Nope only the volume @Abmon98

OpenStudy (abmon98):

is the answer 27

OpenStudy (anonymous):

I have no idea I don't know how to find it haha and it isn't multiple choice

OpenStudy (abmon98):

when both shapes are fitted so they share the same radius and height so 1/3*81

OpenStudy (anonymous):

So then the answer WOULD be 27

OpenStudy (abmon98):

yes

OpenStudy (anonymous):

\[volume~ of~ cylinder~ v=\pi~r ^{2}h\] \[volume~of~cone~ v1=\frac{ 1 }{ 3 }\pi r ^{2}h\] \[\frac{ v1 }{ v }=\frac{ \frac{ 1 }{3 }\pi r ^{2} h}{ \pi r ^{2}h }=\frac{ 1 }{ 3 },v1=\frac{ 1 }{3 }v\]